OpenStudy (anonymous):

Write the scalar equation of the line given the normal vector and point. n= [3,1], P (2,4)

OpenStudy (anonymous):

N = <3i + j>. Let A = <(x-2)i, (y-4)j> be a vector in the line. Then N dot A = 0 because the two vectors are orthogonal. N dot A = 3(x-2) + (y-4) = 0 3x + y = 10 (Answer)

OpenStudy (anonymous):

I don't understand what you did ?